OpenStudy (anonymous):

A 3.3 kg mass with an initial velocity of v1 = aˆi, where a = 3.47 m/s collides with and sticks to a 0.989 kg mass with an initial velocity of v1 = bˆj, where b = −4.8 m/s. Find the final speed of the composite mass. Answer in units of m/s

OpenStudy (anonymous):

I got several answers and they have all been wrong so far :(

OpenStudy (anonymous):

conservation of momentum: \[m_1 v_1 + m_2 v_2 = m_{total} v_{total} = (m_1 + m_2)v_{total}\] \[3.3*(3.47 i) + 0.989*(-4.8 j) = (3.3 + 0.989)*v_{total}\]\[v_{total} = \frac{ 3.3*3.47 }{ 4.289 } i - \frac{ 0.989*4.8 }{ 4.289 }j \]\[\left| v_{total} \right| = \left| \sqrt{v_i^2 + v_j^2} \right|\]\[\theta = \tan^{-1}\left( \frac{ v_j }{ v_i } \right)\] | this is absolute value | lmk if you have questions ^_^

OpenStudy (anonymous):

I got 3.09372 as the v total

OpenStudy (anonymous):

But I inputted it and I got it wrong?

OpenStudy (anonymous):

\[v_i \approx 2.6698\]\[v_j \approx 1.1068\] \[v_tot \approx 2.89\]

OpenStudy (anonymous):

wow thank you :)

OpenStudy (anonymous):

glad i could help ^_^

OpenStudy (anonymous):

could you try another one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

A 7 kg steel ball strikes a wall with a speed of 8.77 m/s at an angle of 59.4 with the normal to the wall. It bounces off with the same speed and angle If the ball is in contact with the wall for 0.237 s, what is the magnitude of the average force exerted on the ball by the wall? Answer in units of N